OpenStudy (anonymous):

someone please help me x=1/ square root (t+1) y=t/(t+1) solve for t x^2= substitute into y y=

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

\[x=\frac{1}{\sqrt{t+1}}\]

OpenStudy (misty1212):

\[\frac{1}{x}=\sqrt{t+1}\\ \frac{1}{x^2}=t+1\\ \frac{1}{x^2}-1=t\]

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

@misty1212 is that the final answer

OpenStudy (misty1212):

it is for "solve for t"

OpenStudy (anonymous):

then substitute into y

OpenStudy (anonymous):

@misty1212